OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, –1) and (–2, –5)? a.5/4 b.-5/4 c.4/5 d.-4/5

OpenStudy (anonymous):

You would have to plug in your two points into m=y2-y1 over x2-x1 So i would be m= -5-(-1) over -2-3 Then Simplify

OpenStudy (anonymous):

so -4/5?

OpenStudy (anonymous):

That is what i got.